May 2, 2001

Jeff Dewynne Mathematical Institute 24–29 St. Giles’ Oxford OX1 3LB, UK +44 1865 270612

Dear Professor Dewynne,

Please consider the enclosed paper, “Inferring model parameters in markets with collars” for publication in Applied Mathematical Finance. Thank you.

Dr. Burton Joseph Rosenberg Dr. Robert Chen burt@cs.miami.edu

(305) 284–2141